3) Solve the area bounded by the curve y = 4x-x and the lines x = -2 and y = 4.

1
Expert's answer
2022-05-25T17:02:45-0400
"4x-x^2=4"




"(x-2)^2=0"

"x=2"


"A=\displaystyle\int_{-2}^{2}(4-(4x-x^2))dy"




"=\big[4x-2x^2+\dfrac{x^3}{3}\big]\begin{matrix}
 2 \\
 -2
\end{matrix}"




"=8-8+\dfrac{8}{3}+8+8+\dfrac{8}{3}=\dfrac{64}{3}({units}^2)"